# Comparative Analysis of Transfer Learning Models for Breast Cancer Classification

### v1

@@ 4.8 Vision Transformer
Table 2 presents a comprehensive overview of the performance outcomes of different deep learning models assessed in this research. It includes precise measurements such as precision, recall, F1-score, and total accuracy metrics for both non-IDC and IDC classifications. This thorough comparison evaluates the performance of the models, offering insights into their efficacy for classifying breast cancer histopathology images, and includes a ranking of their relative performance.
@@
| Resnet-50 | 0.91 | 0.93 | 0.85 | 0.94 | 0.81 | 0.94 | 0.83 | 4.57 (0.051) |
| DenseNet-121 | 0.91 | 0.93 | 0.86 | 0.95 | 0.81 | 0.94 | 0.83 | 4.64 (0.056) |
| ResNeXt-50 | 0.91 | 0.93 | 0.85 | 0.94 | 0.82 | 0.94 | 0.84 | 5.14 (0.055) |
| Inception v3 | 0.91 | 0.93 | 0.84 | 0.94 | 0.84 | 0.94 | 0.84 | 4.57 (0.049) |
| EfficientNet | 0.92 | 0.95 | 0.83 | 0.93 | 0.89 | 0.94 | 0.86 | 3.29 (0.044) |
| MobileNetV2 | 0.90 | 0.94 | 0.82 | 0.93 | 0.85 | 0.93 | 0.83 | 5.71 (0.051) |
| SqueezeNet | 0.88 | 0.95 | 0.74 | 0.88 | 0.87 | 0.91 | 0.80 | 6.21 (0.070) |
| ViT | 0.93 | 0.94 | 0.89 | 0.96 | 0.84 | 0.95 | 0.87 | 1.86 (0.045) |
5 RESULTS AND OBSERVATIONS
5.1 Experimental Setup
5.2 Results and Observations
The ResNet-50 model was trained for a total of 10 epochs, consistently showing a gain in accuracy and a reduction in loss during both the training and validation stages. The model attained a comprehensive accuracy rate of 91%. The non-cancerous forecasts had a precision of 0.93 and a recall of 0.94, while the malignant predictions had a precision of 0.85 and a recall of 0.81. The F1-scores for non-cancerous and malignant classifications were 0.94 and 0.83, respectively. These values demonstrate a balanced precision-recall tradeoff, which is crucial for accurate medical diagnosis.
In the same manner, the DenseNet-121 model underwent training for 10 epochs and attained an overall accuracy of 91%. The accuracy for identifying non-cancerous cases was 0.93, with a sensitivity of 0.95. On the other hand, the accuracy for identifying cancerous cases was 0.86, with a sensitivity of 0.81. The F1-scores for non-cancerous and malignant classifications were 0.94 and 0.83, respectively, indicating the model’s high accuracy in categorizing histopathological pictures.
The ResNeXt-50 model, trained for a duration of 20 epochs, also attained an overall accuracy rate of 91%. The non-cancerous predictions made by the model had a precision of 0.93 and a recall of 0.94, whereas the cancerous predictions had a precision of 0.85 and a recall of 0.82. The F1-scores achieved by this model were 0.94 for non-cancerous classifications and 0.84 for cancerous classifications, indicating its strong performance in classifying medical images.
The GoogLeNet (Inception v3) model, after being trained for 10 epochs, achieved an overall accuracy of 91%. The non-cancerous forecasts had a precision of 0.93 and a recall of 0.94, while the malignant predictions had a precision of 0.84 and a recall of 0.83. The F1-scores for non-cancerous and malignant classifications were 0.94 and 0.84, respectively, demonstrating a robust capacity to maintain a balance between precision and recall.
The EfficientNet model, after undergoing 10 epochs of training, attained the best level of accuracy among the CNN models, reaching 92%. The non-cancerous forecasts achieved a precision of 0.95 and a recall of 0.93, but the malignant predictions had a precision of 0.83 and a recall of 0.89. The F1-scores achieved were 0.94 for non-cancerous classifications and 0.86 for cancerous classifications, indicating that the model exhibited exceptional performance in the field of medical picture classification.
The MobileNetV2 model, trained for 10 epochs, attained a 90% accuracy rate. The benign forecasts exhibited a precision of 0.94 and a recall of 0.93, while the malignant predictions demonstrated a precision of 0.82 and a recall of 0.85. The F1-scores achieved were 0.93 for non-cancerous classifications and 0.83 for cancerous classifications, demonstrating the effectiveness of the system even with limited computational resources.
The SqueezeNet model, after undergoing 10 epochs of training, attained an overall accuracy rate of 88%. The non-cancerous forecasts had a precision of 0.95 and a recall of 0.88, but the cancerous predictions had a precision of 0.74 and a recall of 0.87. The F1-scores for non-cancerous and malignant classifications were 0.91 and 0.80, respectively, indicating its performance in a resource-limited setting.
The Vision Transformer (ViT) model demonstrated superior performance compared to other models. It was trained for 10 epochs and achieved an overall accuracy of 93%. The non-cancerous forecasts achieved a precision of 0.94 and a recall of 0.96, but the malignant predictions had a precision of 0.89 and a recall of 0.84. The F1-scores for the ViT model were 0.95 for non-cancerous classifications and 0.87 for malignant classifications, demonstrating its excellent capacity to successfully balance precision and recall.
5.3 Comparison and Analysis
@@
The Vision Transformer (ViT) model showed improved performance in comparison to conventional CNN designs. ViT achieved a superior overall accuracy of 93%, surpassing other models that had accuracy rates ranging from 88% to 92%. The precision and recall values of ViT for non-cancerous (0.94 and 0.96, respectively) and malignant (0.89 and 0.84, respectively) predictions demonstrate a high degree of dependability in accurately distinguishing between the two classes. The F1-scores of 0.95 for non-cancerous and 0.87 for cancerous classifications highlight the strong ability of the system to properly handle both positive and negative samples.
@@
EfficientNet, one of the CNN models, achieved the maximum accuracy of 92%, demonstrating excellent precision and recall metrics. Nevertheless, it did not quite match the level of performance demonstrated by the ViT model. ResNet, DenseNet, ResNeXt, and GoogLeNet all attained a 91% accuracy, showcasing their efficacy albeit with somewhat inferior precision and recall when compared to EfficientNet and ViT. MobileNetV2 and SqueezeNet, albeit computationally efficient, exhibited lower accuracy (90% and 88% respectively) and somewhat diminished performance metrics, rendering them less effective than the other models for this particular job.
@@
The Vision Transformer (ViT) model is the most successful architecture for classifying invasive ductal carcinoma (IDC) in breast histopathology images. It achieves maximum accuracy and maintains a well-balanced precision, recall, and F1-score. The utilization of self-attention processes in capturing both global and local information inside images provides it with a distinct advantage over conventional CNN architectures.
6 CONCLUSION AND FUTURE WORK

### v2

@@ III-C Results
Table II presents a comprehensive overview of the performance outcomes of different deep learning models assessed in this research. It includes precise measurements such as precision, recall, F1-score, and total accuracy metrics for both non-IDC and IDC classifications. This thorough comparison evaluates the performance of the models, offering insights into their efficacy for classifying breast cancer histopathology images, and includes a ranking of their relative performance.
@@
| Resnet-50 | 0.92 | 0.94 | 0.86 | 0.95 | 0.82 | 0.95 | 0.84 | 4.57 (0.051) |
| DenseNet-121 | 0.92 | 0.94 | 0.87 | 0.96 | 0.82 | 0.95 | 0.84 | 4.64 (0.056) |
| ResNeXt-50 | 0.92 | 0.94 | 0.86 | 0.95 | 0.83 | 0.95 | 0.85 | 5.14 (0.055) |
| Inception v3 | 0.92 | 0.94 | 0.85 | 0.95 | 0.85 | 0.96 | 0.85 | 4.57 (0.049) |
| EfficientNet | 0.93 | 0.96 | 0.84 | 0.94 | 0.90 | 0.95 | 0.87 | 3.29 (0.044) |
| MobileNetV2 | 0.91 | 0.95 | 0.83 | 0.94 | 0.86 | 0.94 | 0.84 | 5.71 (0.051) |
| SqueezeNet | 0.89 | 0.96 | 0.75 | 0.89 | 0.88 | 0.92 | 0.81 | 6.21 (0.070) |
| ViT | 0.94 | 0.95 | 0.90 | 0.97 | 0.85 | 0.96 | 0.88 | 1.86 (0.045) |
III-D Comparison and Analysis
@@
The ViT model showed improved performance compared to conventional CNN designs, achieving a superior overall accuracy of 93%, surpassing other models with accuracy rates ranging from 88% to 92%. The precision and recall values of ViT for non-cancerous (0.94 and 0.96, respectively) and malignant (0.89 and 0.84, respectively) predictions demonstrate a high degree of reliability in accurately distinguishing between the two classes. The F1-scores of 0.95 for non-cancerous and 0.87 for cancerous classifications highlight the system’s strong ability to properly handle both positive and negative samples.
@@
EfficientNet, one of the CNN models, achieved a maximum accuracy of 92%, demonstrating excellent precision and recall metrics. However, it did not quite match the performance level demonstrated by the ViT model. ResNet, DenseNet, ResNeXt, and GoogLeNet all attained a 91% accuracy, showcasing their efficacy, albeit with somewhat inferior precision and recall compared to EfficientNet and ViT. MobileNetV2 and SqueezeNet, although computationally efficient, exhibited lower accuracy (90% and 88%, respectively) and somewhat diminished performance metrics, making them less effective than the other models for this particular task.
@@
The ViT model is the most successful architecture for classifying invasive ductal carcinoma (IDC) in breast histopathology images. It achieves maximum accuracy and maintains a well-balanced precision, recall, and F1-score. The utilization of self-attention mechanisms in capturing both global and local information within images provides it with a distinct advantage over conventional CNN architectures.
IV CONCLUSION AND FUTURE WORK